OpenStudy (anonymous):

how would you find the x intercepts of (x+4)^2-3

OpenStudy (unklerhaukus):

the zeros

OpenStudy (mertsj):

It is not an equation and has no x intercepts.

OpenStudy (anonymous):

that's the thing though, I already have the x's equations as.. sqrt of 3 +- 4?

OpenStudy (unklerhaukus):

\[y(x)=(x+4)^2-3\] x intercept is when y=0 \[0+3=(x+4)^2\] \[√3= ±(x+4)\] \[x=√3-4, √3+4\]

OpenStudy (anonymous):

\[-4\pm \sqrt{3}\]

OpenStudy (anonymous):